Logiweb dictionary of propositional calculus

Up Help

0 0 propositional calculus

1 0 T_{PC} 2 0 A1 3 0 A2 4 0 A3 5 0 MP 6 0 LEMMA1.8 7 2 ∗ ⇒ ∗

The pyk compiler, version 0.grue.20050603 by Klaus Grue GRD-2005-06-06.UTC:10:04:56.802958 = MJD-53527.TAI:10:05:28.802958 = LGT-4624769128802958e-6